# Orbit structures and complexity in Schubert and Richardson Varieties

### v1

@@ 4.2. Orbits in the Schubert Variety
In this section let $w\in W$ and $I\subseteq\Delta$. For any $u\in W$, let $u={{}_{I}u}\prescript{I}{}{u}$ be the left parabolic decomposition of $u$ with respect to $I$. The goal of this section is to use Theorem 4.7 to prove Corollary 1.6 and Theorem 1.5.
Proof of Corollary 1.6. The Bruhat decomposition tells us that the Schubert variety $X_{w}$ is the disjoint union of the Schubert cells for $u\leq w$,
$X_{w}=\displaystyle\bigsqcup_{u\leq w}X_{u}^{\circ}.$
and hence
(6) $\mathcal{O}_{B_{L_{I}}}(X_{w})=\bigsqcup_{u\leq w}\mathcal{O}_{B_{L_{I}}}(X^{\circ}_{u}).$
Theorem 4.7 implies there is a bijection between $\mathcal{O}_{B_{L_{I}}}(X^{\circ}_{u})$ and $\mathcal{O}_{T}(X^{\circ}_{\prescript{I}{}{u}})\times\{u\}$ for all $u\leq w$, which completes our proof. ∎
Remark.
In Corollary 1.6 we decorate each of the sets $\mathcal{O}_{T}(X^{\circ}_{\prescript{I}{}{u}})$ by taking it’s product with $\{u\}$. This is due to the fact that for $u,v\leq w$ with $u\neq v$ it is possible that $\prescript{I}{}{u}=\prescript{I}{}{v}$. And we desire the union in the statement of Corollary 1.6 to be disjoint.
Our next order of business is to apply Corollary 4.8 towards a formula for the minimal codimension of a $B_{L_{I}}$-orbit in any Schubert variety $X_{w}$.
@@
For $w\in W$, the minimal codimension of a $B_{L_{I}}$-orbit in $X_{w}$ is given by
$c_{B_{L_{I}}}(X_{w})=\ell(w)-\max_{u\leq w}\left(\ell(_{I}u)+\operatorname{supp}(\prescript{I}{}{u})\right).$
@@
For a $u\leq w$, the minimal codimension of a $B_{L_{I}}$-orbit in $X^{\circ}_{u}$ is $\ell(\prescript{I}{}{u})-\operatorname{supp}(\prescript{I}{}{u})$ by Corollary 4.8 and Corollary 1.3. The codimension of $X^{\circ}_{u}$ in $X_{w}$ is $\ell(w)-\ell(u)$. These two preceding sentences combine with (6) to imply
$c_{B_{L_{I}}}(X_{w})=$ $\min_{u\leq w}\left(\ell(\prescript{I}{}{u})-\operatorname{supp}(\prescript{I}{}{u})+\ell(w)-\ell(u)\right)$
$=$ $\ell(w)+\min_{u\leq w}\left(-\ell(_{I}u)-\operatorname{supp}(\prescript{I}{}{u})\right)$
$=$ $\ell(w)-\max_{u\leq w}\left(\ell(_{I}u)+\operatorname{supp}(\prescript{I}{}{u})\right),$
where the second inequality is the length-additivity of the left parabolic decomposition. ∎
The following example illustrates that this value is not, in general, well-behaved.
Example 4.10.
Let $w=56781234\in S_{8}$ written in one-line notation. Then $w$ is already in $\prescript{I}{}{W}$ since $\mathcal{D}_{L}(w)\cap I$ is the empty set. Therefore, $\ell(_{I}w)+\operatorname{supp}(\prescript{I}{}{w})=0+7=7$. Now if we apply $s_{4}$ to $w$ and get $u=56718234$, then $\prescript{I}{}{u}=25681347$, and ${}_{I}u$ has length $4$. Therefore, $\ell(_{I}u)+\operatorname{supp}(\prescript{I}{}{u})=4+7=11$. This is not the optimal choice for $u$: if we take $u=53218764\leq w$, then $\ell(_{I}u)+\operatorname{supp}(\prescript{I}{}{u})=6+7=13$.
However, the formula in Proposition 4.9 can be simplified in restricted cases.
Lemma 4.11.
Let $w\in W$ such that $\ell(_{I}w)\geq\ell(_{I}u)$ for all $u\leq w$, i.e. $w$ admits a left Billey-Postnikov decomposition with respect to $I$ (see [RS17, Proposition 4.2]). Then $w$ simultaneously maximizes both $\ell(_{I}u)$ and $\operatorname{supp}(\prescript{I}{}{u})$ among all $u\leq w$. Thus
$c_{B_{L_{I}}}(X_{w})=\ell(\prescript{I}{}{w})-\operatorname{supp}(\prescript{I}{}{w}).$
Proof.
For any $w$, if $u\leq w$, then $\prescript{I}{}{u}\leq\prescript{I}{}{w}$. Thus $w$ always maximizes $\operatorname{supp}(\prescript{I}{}{u})$ among all $u\leq w$. Hence, if $\ell(_{I}w)\geq\ell(_{I}u)$ for all $u\leq w$, then $w$ simultaneously maximizes both $\ell(_{I}u)$ and $\operatorname{supp}(\prescript{I}{}{u})$ among all $u\leq w$. Applying this to Proposition 4.9 yields
$c_{B_{L_{I}}}(X_{w})=$ $\ell(w)-\max_{u\leq w}\left(\ell(_{I}u)+\operatorname{supp}(\prescript{I}{}{u})\right)$
$=$ $\ell(w)-\ell(_{I}w)-\operatorname{supp}(\prescript{I}{}{w})=\ell(\prescript{I}{}{w})-\operatorname{supp}(\prescript{I}{}{w}).$
The hypothesis of Lemma 4.11 is satisfied when $L_{I}$ acts $X_{w}$. This yields one of our main results, Theorem 1.5: a simple, type independent formula for the Levi complexity of a Schubert variety. We are now ready to prove Theorem 1.5.
@@
The equivalence of $L_{I}$ acting on $X_{w}$ and $I\subseteq\mathcal{D}_{L}(w)$ is [BL00, Lemma 8.2.3]. If $I\subseteq{\mathcal{D}}_{L}(w)$, then ${\mathcal{I}}(w)$ contains all the positive roots in the root subsystem generated by $I$. Thus $w=w_{0}(I)d$ is a length-additive expression for some $d\in W$ by [BB05, Proposition 3.1.3]. That is, ${}_{I}w=w_{0}(I)$ in the left parabolic decomposition of $w$. Thus, it must be the case then that $\ell(_{I}w)\geq\ell(_{I}u)$ for all $u\leq w$. Hence Lemma 4.11 implies the theorem statement. ∎

### v2

@@ 4.2. Orbits in the Schubert Variety
In this section let $w\in W$ and $I\subseteq\Delta$. For any $u\in W$, let $u={{}_{I}u}\prescript{I}{}{u}$ be the left parabolic decomposition of $u$ with respect to $I$. Our goal is to extend Corollary 4.8 to a formula for the Levi complexity of a Schubert variety. To do this, we will need to lower bound the codimension of a $B_{L_{I}}$-orbit in $X^{\circ}_{u}$ for $u\leq w$. The added complication in this case is that $L_{I}$ may not act on $X_{u}$, though of course $B_{L_{I}}$ acts, and so $U_{\bar{a}}$ will not be the trivial subgroup of $G$.
For a subvariety $X$ of a variety $Y$, we write $\operatorname{codim}_{Y}(X)$ for the codimension of $X$ in $Y$.
@@
Let $u\leq w\in W$. Let $\Xi\in\mathcal{O}_{B_{L_{I}}}(X^{\circ}_{u})$. Let $h\in\Xi\cap\operatorname{Heart}_{I}(X_{u})$ and $\Theta=T\cdot h$. Then
$\operatorname{codim}_{X^{\circ}_{u}}(\Xi)$ $\geq\ell(u)-\operatorname{supp}(\prescript{I}{}{u})-\ell(w_{0}(I))$
and
$\operatorname{codim}_{X_{w}}(\Xi)\geq\ell(w)-\operatorname{supp}(\prescript{I}{}{u})-\ell(w_{0}(I)).$
@@
The second claimed inequality follows from the first combined with the fact that the codimension of $X^{\circ}_{u}$ in $X_{w}$ is $\ell(w)-\ell(u)$. Hence it remains to prove the first inequality. By [B09, Proposition 1.11] $\dim(\Xi)=\dim(B_{L_{I}})-\dim(B_{h})\leq\dim(B_{L_{I}})-\dim(T_{h})$. Multiplying this inequality by $-1$ and adding $\ell(u)$ yields
(6) $\begin{split}\ell(u)-\dim(\Xi)&\geq\ell(u)-\dim(B_{L_{I}})+\dim(T_{h})\\ &=\ell({{}_{I}u})+\ell(\prescript{I}{}{u})-(\dim(T)+\ell(w_{0}(I)))+\dim(T_{h})\\ &=\ell(\prescript{I}{}{u})-(\dim(T)-\dim(T_{h}))-(\ell(w_{0}(I))-\ell({{}_{I}u}))\\ &=\ell(\prescript{I}{}{u})-\dim(\Theta)-(\ell(w_{0}(I))-\ell({{}_{I}u})),\end{split}$
where the first equality is the length additivity of $u={{}_{I}u}\prescript{I}{}{u}$ and the second equality is [B09, Proposition 1.11]. Proposition 4.2 implies that $\operatorname{Heart}_{I}(X_{u})$ is $T$-equivariantly isomorphic to $X^{\circ}_{\prescript{I}{}{u}}$. Hence, $\dim(\operatorname{Heart}_{I}(X_{u}))=\ell(\prescript{I}{}{u})$ and
(7) $\ell(\prescript{I}{}{u})-\dim(\Theta)=\operatorname{codim}_{\operatorname{Heart}_{I}(X_{u})}(\Theta)\geq c_{T}(X^{\circ}_{\prescript{I}{}{u}})\geq c_{T}(X_{\prescript{I}{}{u}}).$
Thus,
$\operatorname{codim}_{X^{\circ}_{u}}(\Xi)$ $=\dim(X^{\circ}_{u})-\dim(\Xi)=\ell(u)-\dim(\Xi)$
$\geq\ell(\prescript{I}{}{u})-\dim(\Theta)-(\ell(w_{0}(I))-\ell({{}_{I}u}))\geq c_{T}(X_{\prescript{I}{}{u}})-(\ell(w_{0}(I))-\ell({{}_{I}u}))$
$=\ell(\prescript{I}{}{u})-\operatorname{supp}(\prescript{I}{}{u})-(\ell(w_{0}(I))-\ell({{}_{I}u}))=\ell(u)-\operatorname{supp}(\prescript{I}{}{u})-\ell(w_{0}(I)),$
where the first inequality is (6), the second inequality is (7), the third equality is Corollary 1.3, and the final inequality is the length additivity of $u={{}_{I}u}\prescript{I}{}{u}$. ∎
We are now ready to prove Theorem 1.5.
@@
The equivalence of $L_{I}$ acting on $X_{w}$ and $I\subseteq\mathcal{D}_{L}(w)$ is Lemma 4.3.
The Bruhat decomposition tells us that the Schubert variety $X_{w}$ is the disjoint union of the Schubert cells for $u\leq w$, $X_{w}=\bigsqcup_{u\leq w}X_{u}^{\circ}$, and hence
$\mathcal{O}_{B_{L_{I}}}(X_{w})=\bigsqcup_{u\leq w}\mathcal{O}_{B_{L_{I}}}(X^{\circ}_{u}).$
Thus
$c_{B_{L_{I}}}(X_{w})$ $=\min_{\begin{subarray}{c}u\leq w\\ \Xi\in\mathcal{O}_{B_{L_{I}}}(X^{\circ}_{u})\end{subarray}}\operatorname{codim}_{X_{w}}(\Xi)$
$\geq\min_{u\leq w}\left(\ell(w)-\operatorname{supp}(\prescript{I}{}{u})-\ell(w_{0}(I))\right)$
$=\ell(w)-\ell(w_{0}(I))+\min_{u\leq w}\left(-\operatorname{supp}(\prescript{I}{}{u})\right)$
$=\ell(\prescript{I}{}{w})-\max_{u\leq w}\left(\operatorname{supp}(\prescript{I}{}{u})\right)$
$=\ell(\prescript{I}{}{w})-\operatorname{supp}(\prescript{I}{}{w}),$
where the inequality is Proposition 4.9, the third equality follows from Lemma 4.3, and the final equality follows from the fact that $u\leq w$ implies $\prescript{I}{}{u}\leq\prescript{I}{}{w}$.
Corollary 4.8 and Corollary 1.3 imply that this lower bound on $c_{B_{L_{I}}}(X_{w})$ is in fact achieved. We conclude that $c_{B_{L_{I}}}(X_{w})=\ell(\prescript{I}{}{w})+\operatorname{supp}(\prescript{I}{}{w})$. ∎